OpenStudy (anonymous):

write an equation in the form y=mx+b for the line that passes through (1,-10) and has a slope of 3

pooja195 (pooja195):

\[\huge~y+10=3(x-1)\] m=slope (x1,y1) now you can solve to get your answer

OpenStudy (anonymous):

1x3 =3

pooja195 (pooja195):

-1 x 3 =?

OpenStudy (anonymous):

-3

pooja195 (pooja195):

\[\huge~y+10=3x-3\] -3-10 =?

OpenStudy (anonymous):

-13

pooja195 (pooja195):

y=3x-13 <--answer :)

OpenStudy (anonymous):

thank you